OpenStudy (anonymous):

Find the solution to the system of equations. x-4y-2z=1 x+5y+2z=-1 y+3z=10 What about an equation where you have nothing to plug in?

OpenStudy (anonymous):

x-4y-2z=1 x+5y+2z=-1 y+3z=10 (1)---> y=10-3z eliminate 2 equation x-4y-2z=1 x+5y+2z=-1 eliminate x by minus then all follow...... (x-x) + (-4y-5y) + (-2z-2z)= 1-(-1) 0+(-9y) + (-4z) =2 -9y-4z=2 subtitute y with (1) equation -9(10-3z) - 4z =2 -90+27z-4z=2 -90+23z=2 23z=92 z=4 find y,, y=10-3z =10-3(4) =-2 find x by subtitute z and y...choose any equation above.. x-4y-2z=1 x-4(-2)-2(4)=1 x+8-8=1 x=1 .:hence::.x=1,,y=-2,,z=4 prove all by subtitute z,x and y into any equation.... x+5y+2z=-1 (1)+5(-2)+2(4)=-1 1-10+8=-1 -1=-1 proven!!!